OpenStudy (anonymous):

boxes are placed on shelves that are 16 boxes high. if each box is 17 1/2 inches in height, how tall is the stack of boxes?

OpenStudy (anonymous):

17+1/2 = 35/17 16 boxes * 35/17 in/box = 560/17 = 32.94 inches